# Adiabatic Bottlenecks in Quantum Annealing and Nonequilibrium Dynamics of Paramagnons

### v1

@@ Appendix A Gaussian Fluctuations
Appendix B Sherrington-Kirkpatrick Data Set
@@
Below, we present plots contrasting the two data sets by means of the ensemble-averaged exact minigaps (obtained from ARPACK’s largest-magnitude eigenvalues  [^1] Note that, technically, we discard those few instances for which the chosen number of largest-magnitude eigenvalues does not contain the ground and first excited states. Hence, the percentage of hard instances for large $N$ is slightly higher than quoted.), the instantaneous Bogoliubov spectra and the statistical fluctuations. Overall, the magnitude of the fluctuations is found to correlate well with the size of the minigap. The minigap location, however, is not clearly deducible from the aggregate fluctuations $\operatorname{Tr}\bm{F}$. Even so, the fact that the magnitude of the statistical fluctuations correlates so strongly with the size of the minigap can be considered a side result as it confirms the usefulness of the semi-classical approximation to judge the quality of the mean-field solutions. If the bitstring returned by the mean-field is accompanied by small fluctuations, one can be fairly certain that it represents a good solution. This is corroborated by our finding that the adiabatic mean-field dynamics always returns the true ground state when applied to the large-minigap subset of our data. Looking to the left panels of Figs. 10 to 15, this can also be understood to result from the fact that the true minigap is virtually identical to the approximate one given by the Bogoliubov spectrum. Note the heights of the fluctuation peaks increase as we go from Fig. 10 to 15. This also holds for those corresponding to the large-minigap data; however, this is related to the different cut-offs $\Delta_{\mathrm{large}}$ indicated in the right panels of Figs. 16 to 21. In summary, we believe that the mean-field dynamics in combination with the semi-classical fluctuations could be thought of as a “first-order diagnostic” tool that can be used in practice to assess the difficulty of a given problem instance.

### v2

@@ Appendix A Gaussian Fluctuations
Appendix B Maximum 2-Satisfiability
Here we present supplementary results for a hard instance from the maximum-2-satisfiability data set of Ref. [32], where the mean-field method converges to an excited state. Notably, the most frustrated spin in Fig. 9 does not localize at the end of the evolution.
Figure 9: Max. 2-Sat. mean-field trajectories for $N=10$ and $T_{f}=2^{16}$.
This does not render our definition of the localization susceptibility useless, however, as is confirmed by Fig. 10. While we only show one example here to illustrate this point, we have confirmed that these results are robust by considering other instances.
Figure 10: Max. 2-Sat. localization susceptibility for $N=10$ and $T_{f}=2^{16}$.
Appendix C Sherrington-Kirkpatrick Data Set
@@
Below, we present plots contrasting the two data sets by means of the ensemble-averaged exact minigaps (obtained from ARPACK’s largest-magnitude eigenvalues  [^1] Note that, technically, we discard those few instances for which the chosen number of largest-magnitude eigenvalues does not contain the ground and first excited states. Hence, the percentage of hard instances for large $N$ is slightly higher than quoted.), the instantaneous Bogoliubov spectra and the statistical fluctuations. Overall, the magnitude of the fluctuations is found to correlate well with the size of the minigap. The minigap location, however, is not clearly deducible from the aggregate fluctuations $\operatorname{Tr}\bm{F}$. Even so, the fact that the magnitude of the statistical fluctuations correlates so strongly with the size of the minigap can be considered a side result as it confirms the usefulness of the semi-classical approximation to judge the quality of the mean-field solutions. If the bitstring returned by the mean-field is accompanied by small fluctuations, one can be fairly certain that it represents a good solution. This is corroborated by our finding that the adiabatic mean-field dynamics always returns the true ground state when applied to the large-minigap subset of our data. Looking to the left panels of Figs. 12 to 17, this can also be understood to result from the fact that the true minigap is virtually identical to the approximate one given by the Bogoliubov spectrum. Note the heights of the fluctuation peaks increase as we go from Fig. 12 to 17. This also holds for those corresponding to the large-minigap data; however, this is related to the different cut-offs $\Delta_{\mathrm{large}}$ indicated in the right panels of Figs. 18 to 23. In summary, we believe that the mean-field dynamics in combination with the semi-classical fluctuations could be thought of as a “first-order diagnostic” tool that can be used in practice to assess the difficulty of a given problem instance.